find the variance and stand deviatio of the probability distribution of the random variable x, which can take only the values 2,4,5 and 9 given that P(2)=9/10. P(4)=1/20. P(5)= 1/5.P(9)= 3/10

1
Expert's answer
2022-03-04T04:46:32-0500

Check


"9/10+1/20+1/5+3/10=29/20\not=1"

Assume that


"P(2)=9\/20, P(4)=1\/20,"

"P(5)= 1\/5,P(9)= 3\/10"

Check


"9\/20+1\/20+1\/5+3\/10=20\/20=1"


"E(X)=2(9\/20)+4(1\/20)+5(1\/5)+9(3\/10)"

"=4.8"


"E(X^2)=2^2(9\/20)+4^2(1\/20)+5^2(1\/5)+9^2(3\/10)"

"=31.9"

"Var(X)=\sigma^2=E(X^2)-(E(X))^2"

"=31.9-(4.8)^2=8.86"

"\sigma=\sqrt{\sigma^2}=\sqrt{8.86}=2.9766"